If the absorbance of a solution of copper(II) ion decreases by 41.00% upon dilution, what volume of water was added to 20.0 mL of a 1.20 M solution of Cu2+(aq)?
1
Expert's answer
2019-11-18T06:28:48-0500

According to a so-called Beer-Lambert law, the absorbance is linearly proportional to the concentration of the solution:


"A=eCl;"


, where e - absorbance coefficient, C - molar concentration in mol/L and l is a width of the liquid layer.

Thus, we shall try to modify the formula to get the right expression for the concentration of the diluted solution:


"A_1=eC_1l ;"


"A_2=eC_2l;"


"\eta=\frac{A_1-A_2}{A_1}=\frac{el(C_1-C_2)}{elC_1}=\frac{C_1-C_2}{C_1};"


"\eta C_1=C_1-C_2,"


"C_2=C_1(1-\eta),"


"n=C_1V_1,"


"С_2=\frac{n}{V_1+V}=\frac{C_1V_1}{V_1+V}=C_1(1-\eta),"


"\frac{V_1}{V_1+V}=1-\eta,"


Now, we can find a volume of water added:


"V=V_1(\frac{\eta}{1-\eta}),"


Thus, "V=20.0mL(\frac{0.41}{1-0.41})=13.9mL."


There was 13.9 mL of water added.